A 198.4 and a 318.3-kg cart approach each other on a horizontal air track. They collide and stick together. After the collision their total kinetic energy is 1011. What is the speed after the collision?


1
Expert's answer
2022-05-20T04:53:24-0400
"E_k=\frac{mv^2}{2}"

"v=\sqrt{2E_k/m}\\
=\sqrt{2*1011/(198.4+318.3)}=1.978\:\rm m/s"